OpenStudy (didierr):

Jennifer jogged over to Collin's house last Saturday. She first ran from her house to the lake, which took her 15 minutes. She then ran around the lake and over to Collin's house, which took her 24 minutes. Jennifer arrived at Collin's house at 1:14 pm. What time did she leave her own house?

OpenStudy (ddos_dragon):

12:35? 24- 10 = 14 1:14 - 14 = 1:00 10 + 15 = 25 1:00 - 25 minutes 12:35

OpenStudy (ddos_dragon):

Sorry for the late response.

OpenStudy (ddos_dragon):

Remember to medal and close the question. Have a nice day. :)